1. (12pts) Solve the triangle AABC. a. Given A = 45°,b= V8, and c = 2 b. Given a = 10, A = 30°, and b = 24
Answer #1

01. - A- 45° 6=58 C= 2. USE Laws of cosine c 6² +6² -abc son A a² = 4 Je cos (45) a? 12 - 458 x 1 f 92 12-4x257 x 1 11 $4 a =